OpenStudy (anonymous):

What is the exact value of the expression 11√8 + 6√12 − 5√2? Show your work.

OpenStudy (anonymous):

= \[11\sqrt{4.2} + 6\sqrt{4.3} - 5\sqrt{2}\] \[= 22\sqrt{2} + 12\sqrt{3} - 5\sqrt{2}\] \[= 17\sqrt{2} + 12\sqrt{3}\]